# Category $\mathcal{O}$ for Lie superalgebras

### v1

@@ Abstract.
In this paper, the authors define a general Category $\mathcal{O}$ for any quasi-reductive Lie superalgebra. The construction encompasses (i) the parabolic Category ${\mathcal{O}}$ for complex semisimple Lie algebras, and (ii) the known constructions of Category ${\mathcal{O}}$ for specific examples of classical Lie superalgebras. In particular, a parabolic version of Category ${\mathcal{O}}^{\mathfrak{p}}$ for quasi-reductive Lie superalgebras ${\mathfrak{g}}\cong{\mathfrak{g}}_{\overline{0}}\oplus{\mathfrak{g}}_{\overline{1}}$ is developed. In this paper, the authors demonstrate how the parabolic category ${\mathcal{O}}^{{\mathfrak{p}}_{\overline{0}}}$ for the reductive Lie algebra ${\mathfrak{g}}_{\overline{0}}$ plays an important role in the representation theory and cohomology for ${\mathcal{O}}^{\mathfrak{p}}$.
Connections between the categorical cohomology and the relative Lie superalgebra cohomology are firmly established. These results are then used to show that the Category ${\mathcal{O}}$ is standardly stratified. The definition of standardly stratified used in this context is generalized from the original definition of Cline, Parshall, and Scott. Other applications involve showing that the categorical cohomology for ${\mathcal{O}}^{\mathfrak{p}}$ is a finitely generated ring. Furthermore, it is shown that the complexity of modules in Category ${\mathcal{O}}$ is finite with an explicit upper bound given by the dimension of the subspace of the odd degree elements in the given Lie superalgebra.
@@ 1. Introduction
1.1.
The BGG Category $\mathcal{O}$ introduced by Bernstein, I. Gelfand, and S. Gelfand in the early 1970’s has been a central object of study in modern representation theory for nearly fifty years. The BGG Category $\mathcal{O}$ is associated to a complex semisimple Lie algebra $\mathfrak{g}_{{\bar{0}}}$ together with a choice of Borel subalgebra $\mathfrak{b}_{{\bar{0}}}\subseteq\mathfrak{g}_{{\bar{0}}}$. More precisely, it consists of certain possibly infinite-dimensional $U(\mathfrak{g}_{{\bar{0}}})$-modules subject to certain finiteness conditions. Category $\mathcal{O}$ has many remarkable structural properties. For example, it is a highest weight category in the sense of Cline, Parshall and Scott [CPS88], and it admits a Kazhdan–Lusztig theory. Many deep connections to geometry (flag varieties), topology (knot invariants and categorification), as well as combinatorics (via the Kazhdan–Lusztig conjectures) have been established (cf.  [BB81, BK81, EW14, KL79, Str05, Su07]).
Given a parabolic subalgebra $\mathfrak{p}_{{\bar{0}}}$ containing $\mathfrak{b}_{{\bar{0}}}$, one can define a parabolic Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ generalizing Category $\mathcal{O}$. Many of the results for Category $\mathcal{O}$ have analogs in the parabolic setting, in particular Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ is also a highest weight category. Moreover, the representation type of the associated quasi-hereditary algebras was studied in [BN05]. For maximal parabolics in type A, the situation is particularly nice. In this case, these algebras can be constructed diagrammatically via so-called arc algebras (cf.  [BS11, CK14]). Much of the highest weight category structure, such as the Kazhdan–Lusztig polynomials, can be encoded combinatorially in the diagrams (cf.  [BS08]). Slight variations of the arc algebras can then be used to make connections between parabolic Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ and perverse sheaves on Grassmannians [Str09], Springer fibers [SW12], knot homology [Kho02], as well as Lie superalgebras [BS12].
@@
Given the prominence of parabolic Category ${\mathcal{O}}$ for complex semisimple Lie algebras, it is natural to develop the theory of parabolic Category ${\mathcal{O}}$ for quasi-reductive Lie superalgebras ${\mathfrak{g}}={\mathfrak{g}}_{{\bar{0}}}\oplus{\mathfrak{g}}_{{\bar{1}}}$. Category ${\mathcal{O}}$ for Lie superalgebras is not a new concept. It has been developed on a case by case basis (cf. [Br03, Br04, CLW11, CLW12]). For this reason, it is our goal to develop a general theory that provides a unified picture encompassing different versions of Category $\mathcal{O}$ for Lie superalgebras that have appeared in the literature. When ${\mathfrak{g}}={\mathfrak{g}}_{{\bar{0}}}$, this includes the Category $\mathcal{O}$ for a complex semisimple Lie algebra. We remark that some of these categories, e.g. Brundan’s Category $\mathcal{O}_{n}$ [Br04] for the isomeric Lie superalgebras, are not necessarily highest weight categories. Blocks of certain parabolic Category $\mathcal{O}$ which are highest weight categories with finitely many irreducibles have previously been studied in [Maz14, BS24, CCC21].
One of our goals is to highlight the major differences between parabolic Category ${\mathcal{O}}$ for complex semisimple Lie algebras as opposed to parabolic Category ${\mathcal{O}}$ for quasi-reductive Lie superalgebras. For semisimple Lie algebras, the blocks are highest weight categories, and we demonstrate that the general behavior for quasi-reductive Lie superalgebras is that the blocks are standardly stratified categories with possibly infinitely many irreducible representations. In terms of the homological algebra, modules for parabolic Category ${\mathcal{O}}$ for semisimple Lie algebras have finite projective dimension, thus the cohomology lives in finitely many degrees, and have complexity equal to zero. On the other hand, modules for Category ${\mathcal{O}}$ for quasi-reductive Lie superalgebras in general have infinite global dimension and have complexity which can be strictly positive.
@@
From a historical viewpoint, a similar setting via triangular decompositions for finite-dimensional modules was developed by Bellamy and Thiel [BT18]. They generalized the idea of Holmes and Nakano [HN91] by considering the module category over a finite-dimensional algebra $A$ admitting a triangular decomposition. The authors were certainly inspired by the module constructions in [HN91, BT18]. However, many of the arguments cannot be directly applied in our context because of the need to work with infinite-dimensional representations. A more elaborate setup can be found in work of Brundan and Stroppel [BS24].
@@ 1.4. Main Results
    If $M,\ N$ are finite-dimensional modules in $\mathcal{O}^{\mathfrak{p}}$, then $\mathrm{Ext}^{\bullet}_{\mathcal{O}^{\mathfrak{p}}}(M,N)$ is a finitely generated $R$-module.
@@
The finite generation result in Theorem B can be seen as a first step towards developing a support variety theory for our Category $\mathcal{O}^{\mathfrak{p}}$. This would provide a geometric framework that controls the representation theory of $\mathcal{O}^{\mathfrak{p}}$. The authors expect this to be a key tool in studying the complexity of modules.
@@
In Section 2, we formulate a definition of a standardly stratified category that directly generalizes the original definition of a highest weight category due to Cline, Parshall, and Scott via injective modules. Moreover, our definition allows for the category to have infinitely many simples. Using Theorem A, we are able to establish an orthogonality result between standard and costandard type modules. This enables us to prove that our Category $\mathcal{O}^{\mathfrak{p}}$ is standardly stratified in the sense of our new definition.
@@
If $\mathfrak{p}$ is a principal parabolic subalgebra of $\mathfrak{g}$, then the Category $\mathcal{O}^{\mathfrak{p}}$ is standardly stratified in the sense of Definition 2.3.1.
@@
Acknowledgements. We acknowledge Chi-Whi Chen, Shun-Jen Cheng, and Volodymyr Mazorchuk for their input and insights about Category ${\mathcal{O}}$. Many of their comments were incorporated in this paper. Moreover, the authors thank Academia Sinica and the Shenzhen International Center for Mathematics for their hospitality and support during the completion of this paper.
@@ 2. Categories with Filtered Objects
In this section, we present a definition of a standardly stratified category that differs from the one that was first given by Cline, Parshall, and Scott [CPS96]. There are several reasons for stating this definition precisely. First, our definition of standardly stratified is a true generalization of highest weight categories as first defined by Cline, Parshall, and Scott [CPS88]. For example, the categories of rational $G$-modules where $G$ is a reductive algebraic group is a highest weight category. This category has infinitely many simple modules and does not have enough projectives. Thus, this category would not be considered a standardly stratified category under the definitions given in [CPS96] and its variants in [ADL98, D00, Fr07, LW15]. However, with our definition (see Definition 2.3.1), the category of rational $G$-modules is indeed standardly stratified.
@@ 2.2. Highest Weight Categories
For the convenience of the reader, we provide a precise definition of a highest weight category as first given in [CPS88, Definition 3.1]. Recall that a locally Artinian category is a category $\mathcal{C}$ such that any directed union of subobjects lies in $\mathcal{C}$ and every object in $\mathcal{C}$ is a union of subobjects of finite length. The category $\mathcal{C}$ is said to satisfy the Grothendieck condition if
$B\cap(\bigcup A_{i})=\bigcup(B\cap A_{i})\quad\textup{for any subobject }B\textup{ and family }\{A_{i}\}\textup{ of an object }X.$ (2.2.1)
@@
      If $[\nabla(\lambda)/L(\lambda):L(\mu)]\neq 0$, then $\mu<\lambda$.
@@
      For $\lambda,\mu\in\Lambda$, both $\dim\mathrm{Hom}_{\mathcal{C}}(\nabla(\lambda),\nabla(\mu))$ and $[\nabla(\lambda):L(\mu)]$ are finite.
@@ 2.3. Standardly Stratified Categories
Cline, Parshall, and Scott later relaxed the requirements of the highest weight category to introduce standardly stratified categories [CPS96, Definition 2.2.1]. The definitions therein considered only the case when $\Lambda$ is finite (i.e., there are only finitely many simple modules), and used projective covers instead of injective hulls. For our purposes, it makes more sense to extend the definition of standardly stratified categories to include the original definition of the highest weight category. Our presentation differs from other treatments (see [ADL98, D00, Fr07, LW15]), in that (i) one works with injective modules instead of projectives, and (ii) one uses weight posets that could be infinite.
Recall that a quasi-poset (or, a proset) is a set $\Lambda$ having a preorder, i.e., a reflexive transitive relation $\leq$. Elements $i,j\in\Lambda$ are equivalent if and only if $i\leq j$ and $j\leq i$. Denote by $\overline{i}$ the equivalence class of $i$. Thus, $\overline{\Lambda}:=\{\overline{i}~{}|~{}i\in\Lambda\}$ is a poset whose partial order $\leq$ is given by $\overline{i}\leq\overline{j}$ if and only if $i\leq j$.
@@
    A collection $\{\nabla(\lambda)\}_{\lambda\in\Lambda}$ of costandard objects such that if $[\nabla(\lambda):L(\mu)]\neq 0$, then $\mu\leq\lambda$.
@@
      For any $\nu\in\Lambda$, #$\#\{i\geq 0~{}|~{}\mu_{i}=\nu\}<\infty$.
The costandard objects in a standardly stratified category have fewer requirements than those in a highest weight category. For example, if $[\nabla(\lambda)/L(\lambda):L(\mu)]\neq 0$ in a standardly stratified category, then $\mu$ can still be $\lambda$. Moreover, the finiteness conditions (c)(iii) in Definition 2.2.1 on Hom spaces between costandards and lengths of costandards are dropped.

### v2

@@ Abstract.
The authors define a Category $\mathcal{O}$ for any quasi-reductive Lie superalgebra $\mathfrak{g}$ with respect to a triangular decomposition. This much needed approach unifies many important constructions in the existing literature in a rigorous fashion. Our Category $\mathcal{O}$ encompasses all highest weight categories for Lie (super)algebras as well as specific examples which may not be highest weight categories. When the decomposition arises from a principal parabolic subalgebra $\mathfrak{p}$ of $\mathfrak{g}$, the Category $\mathcal{O}$ exhibits rich homological properties. For one, the authors show that in contrast to the case of a semisimple Lie algebra, the Category $\mathcal{O}$ is standardly stratified. Moreover, the categorical cohomology of $\mathcal{O}$ is a finitely generated ring. This provides a first step towards developing a support variety theory for Category $\mathcal{O}$. Finally, it is shown that the complexity of modules in Category $\mathcal{O}$ is finite with an explicit upper bound given by the dimension of the subspace of the odd degree elements in $\mathfrak{g}$. This upgrades results known for $\mathfrak{gl}(m|n)$ to the more general setting. Our arguments are based on foundational connections between the categorical cohomology and the relative Lie superalgebra cohomology as well as the interplay between Category $\mathcal{O}$ for $\mathfrak{g}$ and the Category $\mathcal{O}$ for its corresponding Lie algebra $\mathfrak{g}_{\bar{0}}$.
@@ 1. Introduction
$1.$
The BGG Category $\mathcal{O}$ introduced by Bernstein, I. Gelfand, and S. Gelfand in the early 1970’s has been a central object of study in modern representation theory for nearly fifty years. The BGG Category $\mathcal{O}$ is associated to a complex semisimple Lie algebra $\mathfrak{g}_{{\bar{0}}}$ together with a choice of Borel subalgebra $\mathfrak{b}_{{\bar{0}}}\subseteq\mathfrak{g}_{{\bar{0}}}$. More precisely, it consists of certain possibly infinite-dimensional $U(\mathfrak{g}_{{\bar{0}}})$-modules subject to certain finiteness conditions. Category $\mathcal{O}$ has many remarkable structural properties. For example, it is a highest weight category as defined by Cline, Parshall and Scott [CPS88], and it admits a Kazhdan–Lusztig theory. Deep connections to geometry (flag varieties), topology (knot invariants and categorification), as well as combinatorics (via the Kazhdan–Lusztig conjectures) have been established (cf.  [BB81, BK81, EW14, KL79, Str05, Su07]).
Given a parabolic subalgebra $\mathfrak{p}_{{\bar{0}}}$ containing $\mathfrak{b}_{{\bar{0}}}$, one can define a parabolic Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ generalizing Category $\mathcal{O}$. Many of the results for Category $\mathcal{O}$ have analogs in the parabolic setting, in particular Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ is also a highest weight category. Moreover, the representation type of the associated quasi-hereditary algebras was studied in [BN05]. For other axiomatic approaches via BGG algebras, see Irving [I90].
For maximal parabolics in type A, the situation is particularly nice. In this case, these algebras can be constructed diagrammatically via so-called arc algebras (cf.  [BrS11, CK14]). Much of the highest weight category structure, such as the Kazhdan–Lusztig polynomials, can be encoded combinatorially in the diagrams (cf.  [BrS08]). Slight variations of the arc algebras can then be used to make connections between parabolic Category $\mathcal{O}^{\mathfrak{p}_{{\bar{0}}}}$ and perverse sheaves on Grassmannians [Str09], Springer fibers [SW12], knot homology [Kho02], as well as Lie superalgebras [BrS12].
@@
Given the prominence of parabolic Category $\mathcal{O}$} for complex semisimple Lie algebras, it is natural to develop the theory of parabolic Category $\mathcal{O}$} for quasi-reductive Lie superalgebras ${\mathfrak{g}}={\mathfrak{g}}_{{\bar{0}}}\oplus{\mathfrak{g}}_{{\bar{1}}}$. Category $\mathcal{O}$} for Lie superalgebras is not a new concept. It has been developed on a case by case basis (cf. [Br03, Br04, CLW11, CLW12]), and introduced more generally in [Maz14, AM15].
For this reason, it is our goal to develop a general theory that provides a unified picture encompassing different versions of Category $\mathcal{O}$ for Lie superalgebras that have appeared in the literature. When ${\mathfrak{g}}={\mathfrak{g}}_{{\bar{0}}}$, this includes the Category $\mathcal{O}$ for a complex semisimple Lie algebra. We remark that some of these categories, e.g. Brundan’s Category $\mathcal{O}_{n}$ [Br04] for the isomeric Lie superalgebras, are not necessarily highest weight categories. Blocks of certain parabolic Category $\mathcal{O}$ which are highest weight categories with finitely many irreducibles have previously been studied in [Maz14, CCC21]. One of our goals is to highlight the major differences between parabolic Category $\mathcal{O}$} for complex semisimple Lie algebras as opposed to parabolic Category $\mathcal{O}$} for quasi-reductive Lie superalgebras. For semisimple Lie algebras, the blocks are highest weight categories, and we demonstrate that the general behavior for quasi-reductive Lie superalgebras is that the blocks are standardly stratified categories with possibly infinitely many irreducible representations. In terms of the homological algebra, modules for parabolic Category $\mathcal{O}$} for semisimple Lie algebras have finite projective dimension, thus the cohomology lives in finitely many degrees, and have complexity equal to zero. On the other hand, modules for Category $\mathcal{O}$} for quasi-reductive Lie superalgebras in general have infinite global dimension and have complexity which can be strictly positive.
@@
From a historical viewpoint, a similar setting via triangular decompositions for finite-dimensional modules was developed by Bellamy and Thiel [BT18]. They generalized the ideas of Holmes and Nakano [HN91] by considering the module category over a finite-dimensional algebra $A$ admitting a triangular decomposition. The authors were certainly inspired by the module constructions in [HN91, BT18]. However, many of the arguments cannot be directly applied in our context because of the need to work with infinite-dimensional representations. A more elaborate setup using triangular decompositions can be found in work of Brundan and Stroppel [BrS24].
@@ 1.4. Main Results
    If M,NM,\ N are finite-dimensional modules in $\mathcal{O}^{\mathfrak{p}}$, then $\mathrm{Ext}^{\bullet}_{\mathcal{O}^{\mathfrak{p}}}(M,N)$ is a finitely generated $R$-module.
@@
The finite generation result in Theorem B can be seen as a first step towards developing a support variety theory for our Category $\mathcal{O}^{\mathfrak{p}}$. This would provide a geometric framework for an important piece of the representation theory of $\mathcal{O}^{\mathfrak{p}}$. It is well-known that the cohomology ring is in general too small to control the entire representation theory of $\mathcal{O}^{\mathfrak{p}}$ (cf. [BKN17]), and that a large geometric space, involving detecting subalgebras, will be needed as a key tool in studying the complexity of modules.
@@
In Section 2, we formulate a definition of a standardly stratified category that directly generalizes the original definition of a highest weight category due to Cline, Parshall, and Scott via injective modules (see Theorem 2.4.1). Moreover, our definition allows for the category to have infinitely many simples. Using Theorem A, we are able to establish an orthogonality result between standard and costandard type modules. This enables us to prove that our Category $\mathcal{O}^{\mathfrak{p}}$ is standardly stratified with our new definition.
@@
If $\mathfrak{p}$ is a principal parabolic subalgebra of $\mathfrak{g}$, then the Category $\mathcal{O}^{\mathfrak{p}}$ is standardly stratified as described in Definition 2.3.1.
@@
Acknowledgements. We acknowledge Jonathan Brundan, Chih-Whi Chen, Shun-Jen Cheng, and Volodymyr Mazorchuk for their input and insights about Category $\mathcal{O}$}. Many of their comments were incorporated in this paper. Moreover, the authors thank Academia Sinica, Shenzhen International Center for Mathematics, and the National Center for Theoretical Sciences (NCTS) for their hospitality and support during the completion of this paper.
@@ 2. Categories with Filtered Objects
In this section, we present a definition of a standardly stratified category that differs from the one that was first given by Cline, Parshall, and Scott [CPS96]. There are several reasons for stating this definition precisely. First, our definition of standardly stratified is a true generalization of highest weight categories as first defined by Cline, Parshall, and Scott [CPS88]. For example, the categories of rational $G$-modules where $G$ is a reductive algebraic group is a highest weight category. This category has infinitely many simple modules and does not have enough projectives. Thus, this category would not be considered a standardly stratified category under the definitions given in [CPS96] and its variants in [ADL98, D00, Fr07, LW15, BrS24]. However, with our definition (see Definition 2.3.1), the category of rational $G$-modules is indeed standardly stratified.
@@ 2.2. Highest Weight Categories
For the convenience of the reader, we provide a precise definition of a highest weight category as first given in [CPS88, Definition 3.1]. Recall that a locally Artinian category is a category $\mathcal{C}$ such that any directed union of subobjects lies in $\mathcal{C}$ and every object in $\mathcal{C}$ is a union of subobjects of finite length. The category $\mathcal{C}$ is said to satisfy the Grothendieck condition if, for any subobject $B$ and any family $\{A_{i}\}_{i}$ of subobjects of an object $X$:
$B\cap(\bigcup\nolimits_{i}A_{i})=\bigcup\nolimits_{i}(B\cap A_{i}).$ (2.2.1)
@@
      If $\nabla(\lambda)/L(\lambda):L(\mu)]\neq 0$, then $\mu<\lambda$.
@@
      For $\lambda,\mu\in\Lambda$, both $\dim\mathrm{Hom}_{\mathcal{C}}(\nabla(\lambda),\nabla(\mu))$ and $\nabla(\lambda):L(\mu)]$ are finite.
@@ 2.3. Standardly Stratified Categories
For our purposes, it makes more sense to extend the definition of standardly stratified categories to include the original definition of the highest weight category. Recall that a quasi-poset (or, a proset) is a set $\Lambda$ having a preorder, i.e., a reflexive and transitive relation $\leq$. Elements $i,j\in\Lambda$ are equivalent if and only if $i\leq j$ and $j\leq i$. Denote by $\overline{i}$ the equivalence class of $i$. Thus, $\overline{\Lambda}:=\{\overline{i}~|~i\in\Lambda\}$ is a poset whose partial order $\leq$ is given by $\overline{i}\leq\overline{j}$ if and only if $i\leq j$.
@@
    A collection $\{\nabla(\lambda)\}_{\lambda\in\Lambda}$ of costandard objects such that if $\nabla(\lambda):L(\mu)]\neq 0$, then $\mu\leq\lambda$.
@@
      For any $\nu\in\Lambda$, $\{i\geq 0~|~\mu_{i}=\nu\}<\infty$.
The costandard objects in a standardly stratified category have fewer requirements than those in a highest weight category. For example, if $\nabla(\lambda)/L(\lambda):L(\mu)]\neq 0$ in a standardly stratified category, then $\mu$ can still be $\lambda$. Moreover, the finiteness conditions (c)(iii) in Definition 2.2.1 on Hom spaces between costandards and lengths of costandards are dropped.
2.4.
Originally, Cline, Parshall, and Scott (CPS) severely relaxed the requirements of the highest weight category to introduce standardly stratified categories [CPS96, Definition 2.2.1]. The definitions therein considered only the case when $\Lambda$ is finite (i.e., there are only finitely many simple modules), and used projective covers instead of injective hulls. Moreover, our treatment differs from other approaches (see [ADL98, D00, Fr07, LW15, BrS24]), in that it is not assumed that one is working with a module category for an algebra. With the definition in the prior section, we can present a definitive statement.
Theorem 2.4.1.
If $\mathcal{C}$ is a highest weight category (as defined by CPS in Definition 2.2.1) then $\mathcal{C}$ is standardly stratified as defined in Definition 2.3.1.
Example 2.4.2.
Let $G$ be a reductive algebraic group over an algebraically closed field of characteristic $p>0$. It is well-known the category of rational $G$-modules, $\mathcal{C}=\text{Mod}(G)$, is a prototype for a highest weight category. The category $\mathcal{C}$ has enough injective modules, but does not have enough projective modules. Furthermore, blocks can have infinitely many simple modules.
From Theorem 2.4.1, $\mathcal{C}$ is standardly stratified, however, under older definitions as above, $\mathcal{C}$ would not be viewed as standardly stratified. It is anticipated that Definition 2.3.1 will be useful in the study of the representation theory for quasi-reductive algebraic supergroups.
Example 2.4.3.
The prototype for a standardly stratified category $\mathcal{C}$} is the category of graded modules, $\mathcal{C}=\mathcal{C}_{gr}(A)$ for $A$ a finite-dimensional graded algebra over a field $k$ with a triangular decomposition $A\cong A^{-}\otimes A^{0}\otimes A^{+}$ (as vector spaces).
It was shown in [HN91] and [BT18, Section 4], that $\mathcal{C}$ is standardly stratified under Definition 2.3.1. Again, note that $\mathcal{C}$ can have infinitely many simple modules.
Remark 2.4.4.
From a historical point of view, Cline, Parshall, and Scott (cf. [CPS88]) developed highest weight categories to unify the theory of reductive algebraic group representations in positive characteristic and the Category $\mathcal{O}$} for semisimple Lie algebras over the complex numbers. The authors seek to use terminology that will reflect historical accuracy in the subject.
Highest weight categories encompass the category $A$-fdMod of finite-dimensional modules over a quasi-hereditary finite-dimensional algebra $A$. Our definition of standardly stratified category naturally includes the finite-dimensional algebra setting explained in [CPS96, 2.1, 2.2]. We anticipate that our categorical definition of standardly stratified should encompass ring theoretic framework in [BrS24]. The implications can be described by the diagram below:
[…]
Here, essentially finite fibered quasi-hereditary algebras are defined in [BrS24, Definitions 5.9, 5.18, Sections 2.2–2.4], and $A$-lfdMod is the category of locally finite-dimensional modules over an essentially finite-dimensional locally unital algebra $A$. Examples of $A$-lfdMod that we know are standardly stratified.
It can be deduced from Example 2.4.2 and from [BrS24, Lemma 2.15] that the ring theoretic formulations (on the left hand side) do not include the categorical formulations on the right hand side of the diagram.